OpenStudy (anonymous):

Which of the following expressions allows for the use of the power multiplication formula in order for it to be simplified? ka • kb 2t + 10b w3 ÷ w3 3b • 3n

jhonyy9 (jhonyy9):

ka x kb =k(a+b) 2t +10b w*3 +w*3 =2w*3 (3b)(3n)=3*2(bn)

jhonyy9 (jhonyy9):

the power multiplication can being used when we have one number on exponent on exponent like (2*(3))*2 =2*(6)

jhonyy9 (jhonyy9):

hope that i have helped successfully